OpenStudy (anonymous):

In a game where you can earn positive and negative points, you made these points on the first 5 plays: +10, -42, +39, -25, -5. What is your current number of points in the game?

OpenStudy (anonymous):

Add these numbers up, you can add them in any order. \(+10-42+39-25-5=+10-42+39-30=\) \(10-42+9=19-42=-32\)

OpenStudy (anonymous):

So, here you are looking at just a bunch of adding and subtracting. It might be overwhelming to have 5 numbers but you can just take them one problem at a time and it really isn't too difficult. If you are struggling with adding and subtracting negative numbers, a really good way to get used to the idea is using a number line like the one I posted on your other question. anytime you add a number, move to the right that many times, any time you subtract a number move to the left that many times, hopefully that will help you visualize so the first step is: 10-42 since 42 is a larger number than 10, this answer is going to be a negative number. I find the the best way to approach a situation like this is to subtract 10 from 42 giving you 32 and then making it negative. so, 10-42=-32. now we add that number to the next number so we have: -32+39 which can be rewritten as 39-32 and 39-32=7 now, we need to subtract 25 so: 7-25=-18 now we need to subtract 5 -18-5 Looking at a number line, when you move left 5 spaces from -18 you end up on -23 so, your final answer is: -23